# Evaluating Language Models on Entity Disambiguation in Tables

### v1

@@ 4. Study Set-up
Our evaluation has the main objective of evaluating the selected approaches on datasets not considered in the original experiments, considering especially the evaluation of TURL and TableLlama on STI-derived datasets and of Alligator and Dagobah (with some limitations) on the datasets used to train and evaluate the first two approaches. Therefore, we first discuss the datasets used in our study (Section 4.1), and the evaluation protocol with its associated objectives (Section 4.1). Then, we provide details about ER (Section 4.3) and the training of the models (Section 4.4).
[…]
Gold Standard Dataset Split Tables Cols
min |||| max |||| $\bar{x}$ Rows
min |||| max |||| $\bar{x}$ Entities
Train† 91 1 |||| 8 |||| 4,86 5 |||| 369 |||| 98,61 14 674
2T-red Test† 26 1 |||| 8 |||| 4,65 7 |||| 264 |||| 74,58 4 691
Train 3 691 2 |||| 5 |||| 2,56 4 |||| 8 |||| 5,68 26 189
R1 (HardTables) Test 200 2 |||| 5 |||| 2,59 4 |||| 8 |||| 5,74 1 406
Train† 4 344 2 |||| 5 |||| 2,56 4 |||| 8 |||| 5,57 20 407
SemTab2022 R2 (HardTables) Test 426 2 |||| 5 |||| 2,53 4 |||| 8 |||| 5,56 1 829
Train 9 917 2 |||| 4 |||| 2,51 3 |||| 11 |||| 5,65 64 542
SemTab2023 R1 (WikidataTables) Test 500 2 |||| 4 |||| 2,46 3 |||| 11 |||| 6,95 4 247
120k Train 13 061 1 |||| 43 |||| 5,43 1 |||| 624 |||| 12,98 120 000
TURL 2k-red Test 1 295 1 |||| 14 |||| 1,03 6 |||| 257 |||| 32,95 1 801
@@ 4.1. Datasets
The datasets considered in this work come from different sources and contain information about domains. In particular, we have selected the following:
@@
  • SemTab2022 - R2 (2T) (Cutrona2020, ): the dataset consists of a set of high-quality manually-curated tables with non-obviously linkable cells, i.e., where mentions are ambiguous names, typos, and misspelt entity names;
  • SemTab2022 - R1 & R2 (HardTables) (abdelmageed2022results, ): datasets with tables generated using SPARQL queries (Jimenez2020, ). The datasets used from HardTables 2022 are round 1 (R1) and round 2 (R2). The target KG for this dataset was Wikidata, and as with previous years, the tasks were CEA, CTA, and CPA;
  • SemTab2023 - R1 (WikidataTables) (hassanzadeh2023results, ): datasets with tables generated using SPARQL queries for creating realisticlooking tables. The dataset includes Test and Validation tables, yet we exclusively employ the Validation tables due to Gold Standard (GS) being provided. The target KG for this dataset was Wikidata, and the tasks were CEA, CTA, and CPA;
  • TURL (Deng2022, ): the authors of TURL developed the TURL dataset (Wikitables) using the extensive WikiTable corpus, a rich compilation of tables from Wikipedia. This dataset includes table metadata such as the table name, caption, and column headers. This dataset has been sub-sampled by TableLlama’s authors to create a smaller version containing exactly 2K mentions, called TURL-2K, and used as the test set.
Table 4 reports the statistics of each dataset used for testing (and fine-tuning) (†indicates the datasets that have undergone subsampling). The SemTab datasets were already split in train and test except for BioDiv 2021, which has been only used during the testing phase along with the TURL-2K. To create a unified and common experimental setting, each dataset has been modified to contain the same set of mentions: datasets for TableLlama were first created by generating prompts for each mention while filtering out all prompts that exceeded TableLlama’s context length (which is equal to $8192$ tokens). For this reason, we have renamed BioDiv, 2T and TURL-2K into BioDiv-red, 2T-red and TURL-2K-red. To reduce the number of tokens generated for each prompt, we have taken some precautions: i) the description separator has been reduced from [DESCRIPTION] to [DESC], and ii) the row separator ([SEP]) has been deleted. Then, TURL datasets and the ground truths for Alligator and Dagobah were created using the same mentions as TableLlama.
[…]
Table 3. Characteristics of the datasets used to test the approaches based on ML. For every dataset we report its source, the domain of the entities contained, and the characterization in In-Domain (IN), Out-Of-Domain (OOD) and Moderately-Out-Of-Domain (MOOD) for each of the pre-trained models (see Table 1 for the pre-training datasets). {Dataset}-red means that the specific dataset has been reduced as explained in Section 4.1.
@@
Gold Standard Dataset Source Entity Domain Alligator TURL TableLama
@@
SemTab2022 2T-red Graph Queries / Web Tables Cross IN MOOD MOOD
SemTab2022 R1 (HardTables) Graph Queries Cross IN MOOD MOOD
SemTab2022 R2 (HardTables) Graph Queries Cross IN MOOD MOOD
SemTab2023 R1 (WikidataTables) Graph Queries Cross IN MOOD MOOD
TURL 2K-red Web Tables Cross MOOD IN IN
TURL 120k Web Tables Cross MOOD IN IN
@@ 4.2. Distribution-aware Experimental Objectives
    We assess the performance of the considered approaches on the test data defined in Section 4.1, without further fine-tuning. The heterogeneity of the test data implies all the algorithms are tested against “in-domain”, “moderately-out-of-domain” and “out-of-domain” data, as visible in Table 3
@@
    We fine-tune TURL, TableLlama and Alligator on the train splits from their MOOD data, i.e., on SemTab2022 R1 (HardTables), SemTab2022 R2 (HardTables), SemTab2023 R1 (WikidataTables) and 2T-red for TURL and TableLlama, and on TURL120k [^4] TURL-120k has been created starting from the original TURL dataset (Deng2022, ) by sub-sampling 13’061 tables containing 120’000 mentions ($\approx$ 125’000 is the number of mentions used to fine-tune TableLlama and TURL on their MOOD data combined). for Alligator; finally we test the fine-tuned models on our test data
@@ 4.3. Candidate entity retrieval with LamAPI
The candidates for the mentions contained in the TURL dataset (Deng2022, ), along with its sub-sampled version TURL-2K (Zhang2023, ), were retrieved through the Wikidata-Lookup-Service [^5] https://www.wikidata.org/w/api.php?action=wbsearchentities&search=Obama&language=en&limit=50&format=json, which retrieves at most $50$ candidates for the mention “Obama”, which is known to have a low coverage w.r.t. other ERs (Avogadro2022, ). For this reason, we researched to identify a state-of-the-art approach/tool specific to the ER. The final choice fell on LamAPI, an ER system developed to query and filter entities in a KG by applying complex string-matching algorithms. As suggested in the paper (Avogadro2022, ), we have integrated DBpedia (v. 2016-10 and v. 2022.03.01) and Wikidata (v. 20220708), which are the most popular KGs also adopted in the SemTab challenge666www.cs.ox.ac.uk/isg/challenges/sem-tab. In LamAPI, an ElasticSearch777www.elastic.co index has been constructed, leveraging an engine designed to search and analyse extensive data volumes in nearly real-time swiftly. These customised local copies of the KGs are then used to create endpoints to provide ER services. The advantage is that these services can work on partitions of the original KGs to improve performance by saving time and using fewer resources. This simulates an application setting of large-scale entity disambiguation (large tables), where a local copy can speed up operations substantially. The LamAPI Lookup service was used to extract the candidates, as carried out by other services (Avogadro2023, ). Given a string input, the service retrieves a set of candidate entities from the reference KG.
The choice to use LamAPI as an ER system is based on its availability and performance compared to other available systems (e.g., Wikidata Lookup) (Avogadro2022, ). To further validate the choice of LamAPI we have computed the number of mentions with $K$ candidates for the TURL-2K-red dataset: the original TURL - 2K dataset has almost $600$ mentions with $1$ candidate (the correct one, with $969$ mentions ($\approx 48\%$) with at most 5 candidates included the correct one). For all those mentions the Wikidata-Lookup service fails to retrieve something meaningful. On the contrary LamAPI retrieves for $1650$ mentions ($\approx 91\%$) in our sub-sampled dataset at least $45$ candidates. In particular, for the TURL-2K-red dataset, the coverage (i.e., how many times the correct candidate is retrieved by the ER system over the total number of mentions to cover) of LamAPI and Wikidata-Lookup is $88.17\%$ and $71.75\%$ respectively. For all these reasons we decided to replace the candidates extracted from LamAPI to build a new version of TURL-2K-red dataset which we called TURL-2K-red-LamAPI.
@@ 4.4. Training and implementation details
Pre-training and model usage. For TURL we first replicated the pre-training with the same hyperparameters as specified by the authors in (Deng2022, ) but in a distributed setting on 4 80GB-A100 GPUs, following the findings in (goyal2018accurate, ), then we fine-tuned it with the default hyper-parameters, matching the CEA results in the original paper. We use the open-source version of TableLlama made available on HuggingFace [^8] https://huggingface.co/osunlp/TableLlama. Alligator is pre-trained on different SemTab datasets before 2022 as in the original paper [^9] As reported in Table 3, the 2T dataset is one of the datasets used to pre-train Alligator. Even though we have 2T in our test data, those two datasets come from different rounds and years of the SemTab challenge.. We refer to Table 3 for details about the datasets used for pre-train. Dagobah is run with the default hyperparameters as specified in the corresponding GitHub repository [^10] https://github.com/Orange-OpenSource/Table-Annotation.
@@
Pre-trained Fine-tuned
Dataset TURL TableLlama Alligator Dagobah TURL TableLlama Alligator
@@
BioDiv-red 0,8109 (OOD) 0,9513 (OOD) 0,5674 (OOD) / (OOD) 0,6347 (OOD) 0,9610 (OOD) 0.8547 (OOD)
TURL-2K-red-LamAPI 0,7118 (IN) 0,9051 (IN) 0,6017 (MOOD) o.o.m (MOOD) 0,5347 (IN) 0,9045 (IN) 0,7456 (MOOD $\rightarrow$ IN)
@@
Technical infrastructure. Both test and fine-tuning for TableLlama and TURL has run on a single NVIDIA A100-80GB; Alligator on an AMD EPYC-Milan Processor with 8 cores and 24GB of RAM, while Dagobah on an Intel(R) Xeon(R) CPU E5-2650 0 @ 2.00GHz with $32$ cores and $96$GB of RAM.
@@ 5.1. Main results
Figure 2. Overall elapsed time per dataset.
@@
Overall elapsed time per dataset. Refer to caption
Figure 3. Overall occupied memory per dataset.
[…]
Overall occupied memory per dataset. Refer to caption
Table 4 reports the performances achieved by the four different approaches on the test data both for the pre-trained and fine-tuned models. We observe that Alligator excels on HT-R1, HT-R2 and WikidataTablesR1 SemTab datasets, it performs poorly on both BioDiv-red and TURL-2K-red-LamAPI while performing discretely on 2T-red dataset w.r.t. TableLlama. The opposite is observed for TableLlama and TURL, with TableLlama achieving generally higher accuracy than TURL. This trend can be explained by the fact that HT-R1, HT-R2 and WikidataTables-R1 are considered IN data for Alligator, while both BioDiv and TURL-2K-red-LamAPI, being OOD and MOOD data respectively, contain tables coming from different specific domains, with misspelled, repeated and abbreviated mentions, whose heterogeneity is difficult to capture with handcrafted features. Surprisingly, both TURL and TableLlama excel on BioDiv, even though it’s considered OOD for both approaches: we hypothesise that the enormous and general pre-training knowledge retained by these models explains this excellence. The poor performance achieved by TURL on all the SemTab data, considered MOOD for it, can be explained by the fact that TURL is already fine-tuned on web tables, which are generally coming from a different data distribution than the ones of the SemTab challenge. Interestingly, we observed that both TableLlama and TURL underperforms themselves on the TURL-2K-LamAPI dataset: we argue that the drop in performances, especially for TURL, is due to the increased number of candidates we have retrieved with LamAPI (see the ablation study in section 5.2).
MOOD fine-tuning. Table 4 reports also the performances of TURL, TableLlama and Alligator after the MOOD fine-tuning. Thanks to the “continual finetuning” inspired by (ibrahim2024simple, ) TableLlama slightly increases its performances on (previously) MOOD and OOD (BioDiv-red) data, with no clue of suffering from catastrophic forgetting on its IN data (TURL-2K-red-LamAPI). TURL gains the most from the MOOD fine-tuning ($\approx$$+36\%$ on WikidataTablesR1) but suffers a severe drop in performance on both IN and OOD data, confirming the catastrophic forgetting suffered by models after fine-tuning (goodfellow2015catastrophicforgetting, ). Alligator achieves great performances on 2T-red, BioDiv-red and TURL-2K-red-LamAPI after MOOD fine-tuning, without suffering a huge drop in performance (apart for HT-R2, where it loses 10% points), demonstrating a good generalization pattern.
Computational performance. Figure 2 and Figure 3 report the overall elapsed time for the four tested approaches per dataset and the overall memory occupied by each approach per dataset (TURL and TableLlama’s occupied memory refers to the A100-80GB GPU memory, while the Alligator and Dagobah’s memory refers to the overall RAM of the machine hosting the algorithm). Both time and occupied memory are on a logarithmic scale. If TURL is the fastest and the lightest from both time and occupied memory perspectives, and TableLlama is the slowest and heaviest, Alligator strikes as a good compromise w.r.t. to execution time and occupied memory. Dagobah, on the other hand, is completely out-of-bounds, especially if one is bounded by resource or budget constraints.

### v2

@@ 4. Study Set-up
Our evaluation has the main objective of evaluating the selected approaches on datasets not considered in the original experiments, considering especially the evaluation of TURL and TableLlama on STI-derived datasets and of Alligator and Dagobah (with some limitations) on the datasets used to evaluate the first two approaches. Therefore, we first discuss the datasets used in our study (Section 4.1), and the evaluation protocol with its associated objectives (Section 4.1). Then, we provide details about ER (Section 4.3) and the training of the models (Section 4.4).
@@ 4.1. Datasets
The datasets considered in this work come from different sources and contain information about different domains. In particular, we have selected the following:
@@
  • SemTab2022 - R2 (2T) (Cutrona2020, ): the dataset consists of a set of high-quality manually-curated tables with non-obviously linkable cells, i.e., where mentions are ambiguous names, typos, and misspelled entity names;/
  • SemTab2022 - R1 & R2 (HardTables (HT)) (abdelmageed2022results, ): datasets with tables generated using SPARQL queries (Jimenez2020, ). The datasets used from HardTables 2022 are round 1 (R1) and round 2 (R2);
  • SemTab2023 - R1 (WikidataTables) (hassanzadeh2023results, ): datasets with tables generated using SPARQL queries for creating realisticlooking tables. The dataset includes Test and Validation tables, yet we exclusively employ the Validation tables due to Gold Standard (GS) not being provided;
  • TURL (Deng2022, ): the authors of TURL developed the TURL dataset using the extensive WikiTable corpus, a rich compilation of tables from Wikipedia. This dataset includes table metadata such as the table name, caption, and column headers. This dataset has been sub-sampled by TableLlama’s authors to create a smaller version containing exactly 2K mentions, called TURL-2K, and used as the test set.
Table 3 reports the statistics of each dataset used for testing and fine-tuning the selected approaches (†indicates the datasets that have undergone a reduction). The SemTab datasets were already split in train and test except for BioDiv 2021, which has been only used during the testing phase along with the TURL-2K. To create a unified and common experimental setting, each dataset has been modified to contain the same set of mentions: datasets for TableLlama were first created by generating prompts for each mention while filtering out all prompts that exceeded TableLlama’s context length (which is equal to $8192$ tokens). For this reason, we have renamed BioDiv, 2T and TURL-2K into BioDiv-red, 2T-red and TURL-2K-red. To reduce the number of tokens generated for each prompt, we have taken some precautions: i) the description separator has been reduced from [DESCRIPTION] to [DESC], and ii) the row separator ([SEP]) has been deleted. Then, TURL datasets and the ground truths for Alligator and Dagobah were created using the same mentions as TableLlama.
Table 4. Test datasets denomination, as specified in Section 4.2. For every dataset, we report its source, the domain of the entities contained, and the characterization as In-Domain (IN), Out-Of-Domain (OOD), and Moderately-Out-Of-Domain (MOOD) for each of the pre-trained models (see Table 1 for the pre-training datasets). {Dataset}-red means that the specific dataset has been reduced as explained in Section 4.1.
@@
Dataset Source Entities Domain Alligator TURL TableLlama
SemTab2022 2T-red Graph Queries / Web Tables Cross IN MOOD MOOD
SemTab2022 HT-R1 Graph Queries Cross IN MOOD MOOD
SemTab2022 HT-R2 Graph Queries Cross IN MOOD MOOD
SemTab2023 WikidataTables-R1 Graph Queries Cross IN MOOD MOOD
@@
TURL-120k Web Tables Cross MOOD IN IN
TURL-2K-red Web Tables Cross MOOD IN IN
@@ 4.2. Distribution-aware Experimental Objectives
    We assess the performance of the considered approaches on the test data defined in Section 4.1 and detailed in Table 3, without further fine-tuning. The heterogeneity of the test data implies all the algorithms are tested against “in-domain”, “moderately-out-of-domain” and “out-of-domain” data.
@@
    We fine-tune TURL, TableLlama and Alligator on the train splits from their MOOD data, i.e., on SemTab2022 HT-R1, SemTab2022 HT-R2, SemTab2023 WikidataTables-R1 and SemTab2022 2T-red for TURL and TableLlama, and on TURL120k [^5] TURL-120k has been created starting from the original TURL dataset (Deng2022, ) by sub-sampling 13’061 tables containing 120’000 mentions ($\approx$ 125’000 is the number of mentions used to fine-tune TableLlama and TURL on their MOOD data combined). for Alligator; finally we test the fine-tuned models on our test data.
@@ 4.3. Candidate entity retrieval with LamAPI
The candidates for the mentions contained in the TURL dataset (Deng2022, ), along with its sub-sampled version TURL-2K (Zhang2023, ), were retrieved through the Wikidata-Lookup-Service [^6] https://www.wikidata.org/w/api.php?action=wbsearchentities&search=Obama&language=en&limit=50&format=json, which retrieves at most $50$ candidates for the mention “Obama”, which is known to have a low coverage w.r.t. other ERs (Avogadro2022, ). For this reason, we researched to identify a state-of-the-art approach/tool specific to the ER. The final choice fell on LamAPI, an ER system developed to query and filter entities in a KG by applying complex string-matching algorithms. In LamAPI, an ElasticSearch777www.elastic.co index of the Wikidata dump (v. 01072024) has been constructed, leveraging an engine designed to search and analyse extensive data volumes in nearly real-time swiftly. The choice to use LamAPI as an ER system is based on its availability and performance compared to other available systems (e.g., Wikidata Lookup) (Avogadro2022, ). To further validate the choice of LamAPI we have computed the number of mentions with $K$ candidates for the TURL-2K-red dataset: the original TURL-2K dataset has almost $600$ mentions with $1$ candidate (the correct one, with $969$ mentions ($\approx 48\%$) with at most 5 candidates included the correct one). For all those mentions the Wikidata-Lookup service fails to retrieve something meaningful. On the contrary LamAPI retrieves for $1650$ mentions ($\approx 91\%$) in our sub-sampled dataset at least $45$ candidates. In particular, for the TURL-2K-red dataset, the coverage (i.e., how many times the correct candidate is retrieved by the ER system over the total number of mentions to cover) of LamAPI and Wikidata-Lookup is $88.17\%$ and $71.75\%$ respectively. For all these reasons we decided to replace the candidates extracted from LamAPI to build a new version of TURL-2K-red dataset which we called TURL-2K-red-LamAPI.
@@ 4.4. Training and implementation details
Pre-training and model usage. For TURL we first replicated the pre-training with the same hyperparameters as specified by the authors in (Deng2022, ) but in a distributed setting on 4 80GB-A100 GPUs, following the findings in (goyal2018accurate, ), then we fine-tuned it with the default hyper-parameters, matching the CEA results in the original paper. We use the open-source version of TableLlama made available on HuggingFace [^8] https://huggingface.co/osunlp/TableLlama.
Figure 2. Per-approach average accuracy, elapsed time (min), and occupied memory (GB). The time reported is the elapsed time to link a mention, without considering the ER time. The occupied memory is the occupied GPU memory for TURL and TableLlama, while is the “Virtual Memory Size” for Alligator and Dagobah. {Approach}-FT represents the respective approach fine-tuned in MOOD setting (Sec. 4.2). “O.O.M.” stands for Out-Of-Memory.
@@
Dataset TURL TableLlama Alligator Dagobah GPT-4o GPT-4o-mini TURL-FT TableLlama-FT Alligator-FT
@@
SemTab2021 BioDiv-red 0.8109 OOD 0.9513 OOD 0.5674 OOD N.A. OOD 0.9748 UNK 0.9497 UNK 0.6347 OOD 0.9610 OOD 0.8547 OOD
@@
Alligator is pre-trained on different SemTab datasets before 2022 as in the original paper [^9] As reported in Table 4, the 2T dataset is one of the datasets used to pre-train Alligator. Even though we have 2T in our test data, those two datasets come from different rounds and years of the SemTab challenge.. We refer to Table 1 for details about the datasets used for pre-train. Dagobah is run with the default hyperparameters as specified in the corresponding GitHub repository [^10] https://github.com/Orange-OpenSource/Table-Annotation.
[…]
Technical infrastructure. Both test and fine-tuning for TableLlama and TURL has run on the Azure NC24ads_a100_v4; Alligator on an Intel(R) Xeon(R) Gold 6238R CPU @ 2.20GHz with $64$ cores and $96$GB of RAM, while Dagobah on an Intel(R) Xeon(R) CPU E5-2650 0 @ 2.00GHz with $32$ cores and $96$GB of RAM.
@@ 5.1. Main results
Figure 2 reports the per-approach average accuracy, with the average computed over the test datasets (cfr. Table 3), while Table 5 reports the detailed performances achieved by the four different approaches on the test data. We observe that Alligator excels on HT-R1, HT-R2 and WikidataTablesR1 SemTab datasets, it performs poorly on both BioDiv-red and TURL-2K-red-LamAPI while performing discretely on 2T-red dataset w.r.t. TableLlama. The opposite is observed for TableLlama and TURL, with TableLlama achieving generally higher accuracy than TURL. This trend can be explained by the fact that TURL, TableLlama and Alligator encode different inductive biases: features computed by Alligator aim to capture signals coming from both the underline KG and the syntactical relatedness between a mention and its candidates; GTUM models like TURL and TableLlama on the other hand capture more semantic and high-level features, lacking the direct interaction with the KG. Moreover, the datasets they excel on are the ones considered in-domain (IN). Surprisingly, both TURL and TableLlama excel on BioDiv, even though it’s considered OOD for both approaches: we hypothesise that the enormous and general pre-training knowledge these models retain explains this excellence. The poor performance achieved by TURL on all the SemTab data, considered MOOD for it, can be explained by the fact that TURL is already fine-tuned on web tables, which are generally coming from a different data distribution than the ones of the SemTab challenge. Interestingly, we observed that both TableLlama and TURL underperforms themselves on the TURL-2K-LamAPI dataset: we argue that the drop in performances, especially for TURL, is due to the increased number of candidates we have retrieved with LamAPI (cfr. the ablation study in section 5.2). Regarding the GPT family, GPT-4o is, on average, the best-performing model, with GPT-4o-mini that is on par with the MOOD fine-tuned version of Alligator.
MOOD fine-tuning. Table 5 reports also the performances of TURL, TableLlama and Alligator after MOOD fine-tuning. Thanks to the “continual finetuning” inspired by (ibrahim2024simple, ) TableLlama slightly increases its performances on (previously) MOOD and OOD (BioDiv-red) data, with no clue of suffering from catastrophic forgetting on its IN data (TURL-2K-red-LamAPI). TURL gains the most from the MOOD fine-tuning ($\approx$$+36\%$ on WikidataTables-R1) but suffers a severe drop in performance on both IN and OOD data, confirming the catastrophic forgetting suffered by models after fine-tuning (goodfellow2015catastrophicforgetting, ). Alligator achieves great performances on 2T-red, BioDiv-red and TURL-2K-red-LamAPI after MOOD fine-tuning, while suffering a contained drop in performance (apart from HT-R2, where it loses 10%), demonstrating a better generalization pattern than TURL. The average performance with respect to the data typology is depicted in Figure 3, which further demonstrates these findings.
@@
Figure 3. Per-approach accuracy with respect to data typology, as defined in Section 4.2. The average is computed over datasets sharing the same data typology.
Computational performance. Figure 2 also reports the average elapsed time (center) for the six tested approaches and the average occupied memory (right) by each approach (TURL and TableLlama’s occupied memory refers to the GPU memory, while the Alligator and Dagobah’s memory refers to the overall “Virtual Memory Size”). Note that time is on a logarithmic scale and is considered as the time employed by an approach to link a mention without considering the ER time [^11] The GPT-4o and GPT-4o-mini time is constrained by the limits imposed by OpenAI API rate limits. If TURL is the fastest and the lightest from both time and occupied memory perspectives and TableLlama is the slowest and heaviest, Alligator strikes as a good compromise w.r.t. to execution time and occupied memory. Dagobah, on the other hand, is completely out-of-bounds, especially if one is bounded by resource or budget constraints.